. Select the warmer temperature in each pair (Show the conversions):

a) 10°C or 10°F b) 30°C or 15°F c) -10°C or 32°F d) 200°C or 200°K


1
Expert's answer
2021-05-31T02:26:03-0400

T(°F) = T(°C) × 1.8 + 32

T(K) = T(°C) + 273.15


Solution:

a) 10°C or 10°F

10°C: T(°F) = 10°C × 1.8 + 32 = 50°F

Thus: 10°C warmer than 10°F

Answer (a): 10°C


b) 30°C or 15°F

30°C: T(°F) = 30°C × 1.8 + 32 = 86°F

Thus: 30°C warmer than 15°F

Answer (b): 30°C


c) -10°C or 32°F

-10°C: T(°F) = -10°C × 1.8 + 32 = 14°F

Thus: 32°F warmer than -10°C

Answer (c): 32°F


d) 200°C or 200 K

200°C: T(K) = 200°C + 273.15 = 473.15 K

Thus: 200°C warmer than 200 K

Answer (d): 200°C